OpenStudy (anonymous):

What is the molarity of 60.0g of NaOH in 0.250L of solution? I know that 1mole of NaOH = 40.0g of NaOH... But what I don't understand is how to get to 40g of NaOH?

OpenStudy (chmvijay):

40g 1000 ml gives 1 M 60 g in 250ml gives =1000*60 / 40*250 =6M

OpenStudy (anonymous):

we know that : C=N/V and N=m/M whith C: is the concentration (moles/L) N: nbr of moles ( moles) m : masse ( g ) = 60.0 g M: masse molaire ( NaCl = 58,44 g/mol ) so : C = 4,106776181 mol/L us i know ...........

OpenStudy (preetha):

When ever you are given g of anything, rememeber to convert to moles using: moles = #g /molar mass (g/mole). So to convert 60 g to moles, divide 60 g by 40 g/mol. This # moles is dissolve in .25L. Molarity = # moles /vol in litres. So from the previous calculation, substitute in the numerator, then divide by .25L Or Molarity of NaOH = (60 g/40 g/mol ) x 1/0.25L = ## moles/L = M